OpenStudy (anonymous):

5x-d/6 =2x a. d= 7/6x b. d = 7x c. d = – 7x d. d= -7/6x

OpenStudy (johnweldon1993):

Woops \[\large 5x - \frac{d}{6} = 2x\] Like that?

OpenStudy (anonymous):

um. i think

OpenStudy (johnweldon1993):

Nope I tried it like that and it doesn't work lol

OpenStudy (johnweldon1993):

So I'm actually going to assume it is \[\large \frac{5x - d}{6} = 2x\]

OpenStudy (dmb1215):

what are you trying to do

OpenStudy (anonymous):

solve the equation for D

OpenStudy (johnweldon1993):

So you would multiply both sides by 6 \[\large 5x - d = 12x\] Now subtract 5x from both sides of the equation \[\large -d = 7x\] And finally divide both sides by '-' \[\large d = -7x\]

OpenStudy (dmb1215):

ok

OpenStudy (dmb1215):

wait a sec wouldnt you multiply the 5x to both sides to get rid of 5x by the d

OpenStudy (johnweldon1993):

At what part? @Dmb1215

OpenStudy (dmb1215):

the first part 5x-d/6=2x -d/6=2x*5x

OpenStudy (johnweldon1993):

No...so we have \[\large \frac{5x - d}{6} = 2x\] The first thing you need to do is get rid of that fraction...because you cannot do anything with a fraction on 1 side... That is why I multiplied by 6...it cancels out the fraction... \[\large 5x - d = 12x \] Them it's just combining like terms....

OpenStudy (dmb1215):

ok